OpenStudy (anonymous):

Which set of expressions represents three consecutive integers? A. n, n+1, n+2 B. n, n+2, n+3 C. n, n+2, n+4 D. n, n+3, n+5 Medal and Fan for best/correct answer

OpenStudy (tkhunny):

"Consecutive" This is the key. This is an eyeball problem. How far apart are "consecutive" integers?

OpenStudy (er.mohd.amir):

consecutive integer means integer such that 1,2,3 here u can thick that add one(1) or subtract one(1) get consecutive integers.

OpenStudy (anonymous):

I'm confused about both of your responses... Could you specify it some more?

OpenStudy (tkhunny):

You didn't answer my question. How far apart are "consecutive" integers?

OpenStudy (er.mohd.amir):

if n is a number then for find more cosecutive numbers with n u can add 1 get first consecutive number

OpenStudy (anonymous):

@tkhunny I don't know @Er.Mohd.AMIR So it narrows down to A and B??

OpenStudy (er.mohd.amir):

if number is 5 what is consecutive numbers with 5 that are 5+1,5+2,5+3=6,7,8 are three consecutive numbers with 5

OpenStudy (tkhunny):

If cars pass by your house like this, red, blue, blue, red, yell, silver, green. If you count consecutive cars passing by your house, how many cars will you count?

OpenStudy (er.mohd.amir):

consecutive means "contineous one after one" think about it

OpenStudy (anonymous):

If i'm interpreting this correctly you'd only count the car colors that pass more than once. So the only car colors that passed more than once, were red (2) and blue (2). In total 4 consecutive cars passed by?

OpenStudy (tkhunny):

Why? Who said anything about more than once? If you count consecutive cars, you don't skip any. You get them ALL. There were 7. Consecutive integers are the same. Integers without skipping. 3, 4, 5, 6, 7 8, 9, 10, 11, 12 -1, -2, -3, -4 Nonconsecutive 1, 2, 5, 6, 10 3, 5, 7, 9 2, 4, 8, 16 Consecutive integers move by just one when they move. Nonconsecutive can move by more. Choice D n, n+3, n+5 The first move is from n to n+3. This is a jump of 3. That is not a jump of 1. This is nonconsecutive. Choice C. n, n+2, n+4 The first move is from n to n+2. This is a jump of 2. That is not a jump of 1. This is nonconsecutive. You judge Choices B and A.

OpenStudy (anonymous):

Oh, So both a and b are correct or is there something else to do?

OpenStudy (er.mohd.amir):

if u write numbers in math like 1,2,3,4.................... what u do here just add 1 in previous number 1,1+1,2+1,3+1,............................ these numbers are consecutive.

OpenStudy (anonymous):

So A?

OpenStudy (er.mohd.amir):

yes

OpenStudy (anonymous):

Thank you @Er.Mohd.AMIR

OpenStudy (tkhunny):

You can also subtract and check. Are these consecutive? 1, 3, 4, 5, 6 6-5 = 1 -- Maybe 5-4 = 1 -- Maybe 4-3 = 1 -- Maybe 3-1 = 2 -- Whoops! Not Consecutive.